OpenStudy (anonymous):

At a local sawmill sawdust is falling onto a pile at a rate of 0.5pi m^3/min.If the pile maintains the shape of a cone with the height equal to three times the radius of the base,how fast is the height of the pile increasing when the volume is 8pi.m^3?

OpenStudy (anonymous):

Help!!

OpenStudy (mertsj):

\[V=\frac{1}{3}\pi r^2h\] \[h=3r, r=\frac{h}{3}\] \[V=\frac{1}{3}\pi (\frac{h}{3})^2h=\frac{1}{27}\pi h^3\] When Volume = 8pi \[8\pi = \frac{1}{27} \pi h^3\] \[8(27)=h^3\] \[h=6\] \[\frac{dV}{dt}=\frac{1}{9}\pi h^2\frac{dh}{dt}\] \[.5\pi=\frac{1}{9}\pi (6^2)\frac{dh}{dt}\] \[.5\pi =4\pi \frac{dh}{dt}\] \[\frac{dh}{dt}=\frac{1}{8}\]